OpenStudy (anonymous):

Condense the expression. log714 - log75

OpenStudy (anonymous):

Log(714/75)

OpenStudy (mathmagician):

log714-log75=log(714/75)=log(238/25)

OpenStudy (anonymous):

i ment Log7 14 - Log7 5

OpenStudy (anonymous):

Log7(14/5)

OpenStudy (anonymous):

\[1+\log_{7}\left(\frac{2}{5}\right)\]

OpenStudy (anonymous):

This question really depends on what "condensed form" is.